# Car-1000: A New Large Scale Fine-Grained Visual Categorization Dataset

### v1

@@ Abstract
Fine-grained visual categorization (FGVC) is a challenging but significant task in computer vision, which aims to recognize different sub-categories of birds, cars, airplanes, etc. Among them, recognizing models of different cars has significant application value in autonomous driving, traffic surveillance and scene understanding, which has received considerable attention in the past few years. However, Stanford-Car, the most widely used fine-grained dataset for car recognition, only has 196 different categories and only includes vehicle models produced earlier than 2013. Due to the rapid advancements in the automotive industry during recent years, the appearances of various car models have become increasingly intricate and sophisticated. Consequently, the previous Stanford-Car dataset fails to capture this evolving landscape and cannot satisfy the requirements of automotive industry. To address these challenges, in our paper, we introduce Car-1000, a large-scale dataset designed specifically for fine-grained visual categorization of diverse car models. Car-1000 encompasses vehicles from 165 different automakers, spanning a wide range of 1000 distinct car models. Additionally, we have reproduced several state-of-the-art FGVC methods on the Car-1000 dataset, establishing a new benchmark for research in this field. We hope that our work will offer a fresh perspective for future FGVC researchers. Our dataset is available at https://github.com/toggle1995/Car-1000.
@@ 1 Introduction
To address the aforementioned problems, in this paper, we introduce Car-1000, a novel and large-scale comprehensive dataset for fine-grained car classification. First, to ensure the dataset aligns with the latest developments in the automotive industry, we extract the popularity and user comments of different car models from one of the world’s largest automotive forums, “DongCheDi” [1], and select 1000 high-attention categories to be included in our dataset. Then, we acquire images of each category from the internet through web scraping scripts and employ professional annotators with considerable knowledge about cars to meticulously filter the images. The whole process cost us more than 4000 dollars. Consequently, we present a new, large-scale fine-grained car dataset containing 1000 categories and 140312 images. The whole dataset has been released to the community, thus facilitating future research endeavors in this field.
@@
    Our Car-1000 dataset is a large-scale comprehensive dataset, comprising 140312 images across 1000 different models from 165 automakers. Considering both the total number of classes, the volume of images, and the diversity of the included automakers, our Car-1000 emerges as the largest and most comprehensive dataset towards fine-grained car classification.
@@ 2.2 Dataset Characteristics
Relative Large-scale dataset. Our Car-1000 encompasses 140312 images, spanning 1000 categories from 165 automakers. Considering the number of automakers, images, classes and the average images per class, our Car-1000 stands out as the most comprehensive dataset, which surpasses the previous ones such as Stanford Cars [19], CompCars [28], BVMMR [2].
@@ 4 Conclusion
In this paper, we introduce Car-1000, a new large-scale comprehensive fine-grained car classification dataset. Compared to the existing datasets, our Car-1000 stands out for the following advantages. First, Car-1000 is one of the largest datasets dedicated to car classification with 138411 images, covering 1000 different models. It is worthwhile to mention that these 1000 different models come from 165 different automakers, endowing Car-1000 with a high level of diversity. Second, Car-1000 exhibits extensive temporal coverage, including the models from 1960s to the latest 2020s. By doing so, Car-1000 dataset is the most up-to-date one in this field, which captures the new trend of car industry. Third, Car-1000 incorporates a three-tier hierarchical labeling system with 7 primary-categories, 21 secondary-categories and 1000 sub-classes, which provides the attribute and size information of each involved model. Additionally, we benchmark several state-of-the-art classification networks on Car-1000, which serves as the baseline results for future researchers. We hope our work will provide a new perspective for future research of FGVC.

### v2

@@ Abstract
Fine-grained visual categorization (FGVC) is a challenging but significant task in computer vision, which aims to recognize different sub-categories of birds, cars, airplanes, etc. Among them, recognizing models of different cars has significant application value in autonomous driving, traffic surveillance and scene understanding, which has received considerable attention in the past few years. However, Stanford-Car, the most widely used fine-grained dataset for car recognition, only has 196 different categories and only includes vehicle models produced earlier than 2013. Due to the rapid advancements in the automotive industry during recent years, the appearances of various car models have become increasingly intricate and sophisticated. Consequently, the previous Stanford-Car dataset fails to capture this evolving landscape and cannot satisfy the requirements of automotive industry. To address these challenges, in our paper, we introduce Car-1000, a large-scale dataset designed specifically for fine-grained visual categorization of diverse car models. Car-1000 encompasses vehicles from 166 different automakers, spanning a wide range of 1000 distinct car models. Additionally, we have reproduced several state-of-the-art FGVC methods on the Car-1000 dataset, establishing a new benchmark for research in this field. We hope that our work will offer a fresh perspective for future FGVC researchers. Our dataset is available at https://github.com/toggle1995/Car-1000.
@@ 1 Introduction
To address the aforementioned problems, in this paper, we introduce Car-1000, a novel and large-scale comprehensive dataset for fine-grained car classification. First, to ensure the dataset aligns with the latest developments in the automotive industry, we extract the popularity and user comments of different car models from one of the world’s largest automotive forums, “DongCheDi” [7], and select 1000 high-attention categories to be included in our dataset. Then, we acquire images of each category from the internet through web scraping scripts and employ professional annotators with considerable knowledge about cars to meticulously filter the images. The whole process cost us more than 4000 dollars. Consequently, we present a new, large-scale fine-grained car dataset containing 1000 categories and 140267 images. The whole dataset has been released to the community, thus facilitating future research endeavors in this field.
@@
    Our Car-1000 dataset is a large-scale comprehensive dataset, comprising 140267 images across 1000 different models from 166 automakers. Considering both the total number of classes, the volume of images, and the diversity of the included automakers, our Car-1000 emerges as the largest and most comprehensive dataset towards fine-grained car classification.
@@ 2.2 Dataset Characteristics
Relative Large-scale dataset. Our Car-1000 encompasses 140267 images, spanning 1000 categories from 166 automakers. Considering the number of automakers, images, classes and the average images per class, our Car-1000 stands out as the most comprehensive dataset, which surpasses the previous ones such as Stanford Cars [19], CompCars [28], BVMMR [1].
@@ 4 Conclusion
In this paper, we introduce Car-1000, a new large-scale comprehensive fine-grained car classification dataset. Compared to the existing datasets, our Car-1000 stands out for the following advantages. First, Car-1000 is one of the largest datasets dedicated to car classification with 140267 images, covering 1000 different models. It is worthwhile to mention that these 1000 different models come from 166 different automakers, endowing Car-1000 with a high level of diversity. Second, Car-1000 exhibits extensive temporal coverage, including the models from 1960s to the latest 2020s. By doing so, Car-1000 dataset is the most up-to-date one in this field, which captures the new trend of car industry. Third, Car-1000 incorporates a three-tier hierarchical labeling system with 7 primary-categories, 21 secondary-categories and 1000 sub-classes, which provides the attribute and size information of each involved model. Additionally, we benchmark several state-of-the-art classification networks on Car-1000, which serves as the baseline results for future researchers. We hope our work will provide a new perspective for future research of FGVC.